# XAMPLER: Learning to Retrieve Cross-Lingual In-Context Examples

### v2

@@ Abstract
Recent studies indicate that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving relevant in-context examples tailored to the input query, enhances few-shot in-context learning of English. However, adapting these methods to other languages, especially low-resource ones, poses challenges due to the scarcity of cross-lingual retrievers and annotated data. Thus, we introduce XAMPLER: Cross-Lingual Example Retrieval, a method tailored to tackle the challenge of cross-lingual in-context learning using only annotated English data. XAMPLER first trains a retriever based on Glot500, a multilingual small language model, using positive and negative English examples constructed from the predictions of a multilingual large language model, i.e., MaLA500. Leveraging the cross-lingual capacity of the retriever, it can directly retrieve English examples as few-shot examples for in-context learning of target languages. Experiments on the multilingual text classification benchmark SIB200 with 176 languages show that XAMPLER substantially improves the in-context learning performance across languages. Our code is available at https://github.com/cisnlp/XAMPLER.
@@ 1 Introduction
Recent studies (Luo et al., 2024) have uncovered a more strategic approach to example retrieval. Rather than relying on random selection, these studies advocate for retrieving examples tailored to the input query, resulting in notable performance enhancements in in-context learning. The retrievers employed by these methods can be categorized into two main types: general off-the-shelf retrievers (Liu et al., 2022), e.g., Sentence-BERT (Reimers and Gurevych, 2019), and task-specific fine-tuned retrievers (Rubin et al., 2022), which are trained based on LLM signals using labeled data.
Figure 1: XAMPLER involves three steps: 1. Data Construction: given a query in English $q_{i}$, we divide the candidate examples $D_{i}$ into positive examples $D_{i}^{pos}$ and negative examples $D_{i}^{neg}$ based on the prediction of MaLA500 (Lin et al., 2024b); 2. Retriever Fine-tuning: we fine-tune the retriever based on Glot500 (Imani et al., 2023) using the constructed data; 3. In-Context Learning: given a query in any language $q_{j}$, we use the fine-tuned retriever to retrieve relevant English examples as few-shots for in-context learning. During training, XAMPLER is English-only, but for evaluation via in-context learning, it extends to any of the 500+ languages supported in MaLA500 and Glot500.
@@
To tackle these challenges, we propose a simple yet effective method that relies solely on annotated English data, termed XAMPLER (Cross-Lingual Example Retrieval). As shown in Fig. 1, given an English query $q_{i}$ and an English example from the candidate pool $D_{i}$, we employ in-context learning with MaLA500 (Lin et al., 2024b), a 10B multilingual LLM covering 534 languages, to predict the label of the query. Based on the correctness of the prediction, we classify the candidate example as either positive or negative, i.e., $D_{i}^{pos}$ and $D_{i}^{neg}$. Then, leveraging the curated dataset, we train a retriever based on Glot500 (Imani et al., 2023), a multilingual small language model covering 534 languages, aiming to minimize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023). Finally, the trained retriever is directly applied to retrieve valuable few-shot examples in English for the given query in the target language. The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for in-context learning. Experiments across 176 languages on SIB200 show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning across languages.
@@ 2.2 Data Construction
Scoring all pairs of training examples presents a quadratic complexity in $|D|$, making it resource-intensive. Inspired by Rubin et al. (2022), we mitigate this by selecting the top $k$ similar examples as candidates. We utilize Sentence-BERT (SBERT) (Reimers and Gurevych, 2020) [^1] We use version distiluse-base-multilingual-cased-v1. for candidate selection. Based on our experiments detailed in Section C, we set $k=10$. The top $k$ candidates for $q_{i}$ are denoted as $D_{i}=\{d_{i,1},\cdots,d_{i,k}\}$, where each candidate $d_{i,j}$ is represented as $(x_{i,j},y_{i,j})$, with $x_{i,j}$ being the input and $y_{i,j}$ the corresponding label.
@@ 2.3 Retriever Fine-tuning
We utilize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023) to train the task-specific retriever, aiming to maximize the similarity between $q_{i}$ and $x_{i,j}$ if $x_{i,j}$ is a positive example while minimizing the similarity if $x_{i,j}$ is a negative example. We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant cost of fine-tuning an LLM. We train for 50 epochs with a learning rate of 2e-5. Due to the multilingual nature of Glot500, the fine-tuned retriever can be effectively transferred to retrieve in-context examples for other languages.
@@ 3.1 Setup
We evaluate XAMPLER on a massively multilingual text classification benchmark, SIB200 (Adelani et al., 2023). SIB200 involves seven classes: science/technology, travel, politics, sports, health, entertainment, and geography. Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500 (see §A). The English training set contains 701 samples, and each language has 204 samples for evaluation.
Our evaluation framework follows the prompt template used in Lin et al. (2024b): ‘The topic of the news [sent] is [label]’, where [sentence] represents the text for classification and [label] is the ground truth. [label] is included when the sample serves as a few-shot example but is omitted when predicting the sample. We opt for English prompt templates over in-language ones due to the labor-intensive nature of crafting templates for non-English languages, especially those with limited resources. MaLA500 takes the concatenation of few-shot examples and $q_{i}$ as input, then proceeds to estimate the probability distribution across the label set. We measure the performance with accuracy.
@@
Off-the-shelf Retriever. We utilize SBERT (Reimers and Gurevych, 2019), a model covering 50+ languages trained with parallel corpora based on mBERT (Devlin et al., 2019). Additionally, we employ two massively multilingual language models, namely Glot500 and MaLA500, as retrievers, denoted as Glot500 RET and MaLA500 RET, respectively. Tailored examples are retrieved based on the cosine similarity between the sentence representations of the candidate and the query. For Glot500, we utilize mean pooling over hidden states of the selected layer. For MaLA500, we adopt a position-weighted mean pooling method on the selected layer, assigning higher weights to later tokens (Muennighoff, 2022). We use K-Nearest Neighbors (KNN) to select the layer that performs best across layers (see §B). The selected layers for Glot500 and MaLA500 are 11 and 21, respectively.
Cross-lingual Transfer. Cross-lingual transfer is another baseline exploiting English data. In this approach, the multilingual language model is fine-tuned with English data and then deployed for evaluation across target languages. Both Glot500 and MaLA500 are included, with their corresponding cross-lingual transfer baselines denoted as Glot500 XLT and MaLA500 XLT. For Glot500, we opt for full-parameter fine-tuning. For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we only update the LoRA parameters with prompt tuning.
K-Nearest Neighbors. We consider K-Nearest Neighbors (KNN) with the fine-tuned task-specific retriever of XAMPLER and the baselines with retrievers for comparison. Specifically, we adopt majority voting based on the labels of the examples retrieved by the given retriever.
KNN ICL
Latin Non-Latin Avg Latin Non-Latin Avg
Random - - - 56.38 58.66 57.14
SBERT 44.90 37.13 42.29 63.90 62.93 63.57
Glot500 RET 51.16 60.26 54.21 65.21 70.09 66.85
MaLA500 RET 31.90 32.63 32.15 61.02 63.58 61.88
Glot500 XLT - - - 67.09 74.30 69.51
MaLA500 XLT - - - 69.15 71.39 69.90
XAMPLER 67.05 75.97 70.04 73.59 79.80 75.67
Table 1: Average macro-accuracy across all 176 languages, 117 languages in Latin scripts and 59 in non-Latin scripts on SIB200 using XAMPLER and the baselines. KNN results are only provided for retriever-based methods. Results are based on the 3-shot setting.
@@ 3.2 Main Results
The comparison between the baselines and XAMPLER is illustrated in Table 1. Our analysis reveals several insights based on the performance with In-Context Learning (ICL) across different methods. Notably, the random baseline exhibits the worst performance among the baselines using ICL, emphasizing the critical role of example selection for effective in-context learning. Leveraging an off-the-shelf retriever notably boosts performance. For instance, Glot500 RET outperforms the random baseline by 9.71%, and it also outperforms MaLA500 RET, showcasing the validity of selecting Glot500 as the base retriever. Leveraging English data further enhances performance, with both Glot500 XLT and MaLA500 XLT surpassing their corresponding RET baselines by 2.66% and 8.02%.
Among all the methods, XAMPLER achieves the highest performance, surpassing the second-best method, MaLA500 XLT, by 5.77%. This highlights the effectiveness of training a task-specific retriever solely with English data. Moreover, XAMPLER outperforms MaLA500 XLT by 4.44% on languages written in Latin scripts and by 8.41% on languages written in non-Latin scripts. It shows that XAMPLER can provide greater benefits for languages with non-Latin scripts, which are relatively isolated from English written in Latin.
3.3 Effect of Task-Specific Retriever
The results from various retrievers utilizing KNN, as shown in Table 1, indicate that XAMPLER’s fine-tuned retriever excels at retrieving more examples within the same classes as the query in the target language. Notably, XAMPLER with KNN outperforms the second-best retriever, Glot500 RET, by a notable margin, i.e., 15.83%. This superiority enables XAMPLER to leverage majority label bias in in-context learning (Zhao et al., 2021), thereby enhancing overall performance.
3.4 Effect of In-Context Learning
In Table 1, we compare XAMPLER with KNN against that with ICL. Notably, XAMPLER with ICL surpasses XAMPLER with KNN by 5.63%. Moreover, XAMPLER with ICL demonstrates an average improvement of 6.54% for languages in Latin and 3.83% for languages in non-Latin. This disparity may be attributed to the ability of in-context learning to effectively model queries written in the same script as the provided examples, while facing challenges in handling queries written in different scripts from the few-shot examples.
@@ 3.4 Effect of In-Context Learning
We further compare XAMPLER’s performance with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 2. Interestingly, XAMPLER with ICL exhibits inconsistent superiority over KNN, with performance variances ranging from 3% to 10%. Specifically, XAMPLER with KNN achieves its peak performance with 5 examples, whereas ICL achieves impressive results with only 2 examples. Notably, in comparison to KNN’s optimal performance, recorded at 73.26% with 5 shots, XAMPLER with ICL demonstrates a notable improvement of 2.58%. These findings underscore the efficacy of applying in-context learning in effectively leveraging the retrieved examples.
@@ 4 Related Work
Several works (Shi et al., 2022; Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023) extend these methods to non-English languages. A study closely related to ours is Shi et al. (2022), which trains a cross-lingual example retriever via distilling the LLM’s scoring function and evaluates it on four languages for the Text-to-SQL Semantic Parsing task. However, our contribution lies in addressing the more challenging low-resource scenario, thereby extending the applicability and robustness of the approach proposed by Shi et al. (2022).
@@ 5 Conclusion
In this paper, we introduce XAMPLER, a novel approach designed for cross-lingual example retrieval to facilitate in-context learning in any language. Relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language. Our experiments on SIB200 across 176 languages show that XAMPLER can outperform previous methods by a notable margin.

### v3

@@ Abstract
Recent studies indicate that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving relevant in-context examples tailored to the input query, enhances few-shot in-context learning for English. However, adapting these methods to other languages, especially low-resource ones, poses challenges due to the scarcity of cross-lingual retrievers and annotated data. Thus, we introduce XAMPLER: Cross-Lingual Example Retrieval, a method tailored to tackle the challenge of cross-lingual in-context learning using only annotated English data. XAMPLER first trains a retriever based on Glot500, a multilingual small language model, using positive and negative English examples constructed from the predictions of a multilingual large language model, i.e., MaLA500. Leveraging the cross-lingual capacity of the retriever, it can directly retrieve English examples as few-shot examples for in-context learning of target languages. Experiments on two multilingual text classification benchmarks, namely SIB200 with 176 languages and MasakhaNEWS with 16 languages, demonstrate that XAMPLER substantially improves the in-context learning performance across languages. Our code is available at https://github.com/cisnlp/XAMPLER.
@@ 1 Introduction
Recent studies (Luo et al., 2024) have uncovered a more strategic approach to example retrieval. Rather than relying on random selection, these studies advocate for retrieving examples tailored to the input query, resulting in notable performance enhancements in in-context learning. The retrievers employed by these methods can be categorized into two main types: general off-the-shelf retrievers (Liu et al., 2022), e.g., Sentence-BERT (Reimers and Gurevych, 2019), and task-specific fine-tuned retrievers (Rubin et al., 2022), which are trained based on LLM signals (whether an example is helpful) using labeled data.
Figure 1: XAMPLER involves three steps: 1. Data Construction: given a query in English $q_{i}$, we divide the candidate English examples $D_{i}$ into positive examples $D_{i}^{pos}$ and negative examples $D_{i}^{neg}$ based on the prediction of MaLA500 (Lin et al., 2024b); 2. Retriever Fine-tuning: we fine-tune the retriever based on Glot500 (Imani et al., 2023) using the constructed data; 3. In-Context Learning: given a query in any language $q_{j}$, we use the fine-tuned retriever to retrieve relevant English examples $D_{j}$ as few-shots for in-context learning. For training, XAMPLER requires English data only. Once trained, the model can be applied to any of the 500 languages covered by MaLA500/Glot500 without any need for (often unavailable) labeled low-resource data.
@@
To tackle these challenges, we propose a simple yet effective method that relies solely on annotated English data, termed XAMPLER (Cross-Lingual Example Retrieval). As shown in Fig. 1, given an English query $q_{i}$ and an English example from the candidate pool $D_{i}$, we employ in-context learning with MaLA500 (Lin et al., 2024b), a 10B multilingual LLM covering 534 languages, to predict the label of the query. Based on the correctness of the prediction, we classify the candidate example as either positive or negative, i.e., $D_{i}^{pos}$ and $D_{i}^{neg}$. Then, leveraging the curated dataset, we train a retriever based on Glot500 (Imani et al., 2023), a multilingual small language model covering 534 languages, aiming to minimize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023). Finally, the trained retriever is directly applied to retrieve valuable few-shot examples in English for the given query in the target language. The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for in-context learning. Experiments across 176 languages on SIB200 and 16 languages on MasakhaNEWS show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning across languages.
@@ 2.2 Data Construction
Scoring all pairs of training examples presents a quadratic complexity in $|D|$, making it resource-intensive. Inspired by Rubin et al. (2022), we mitigate this by selecting the top $k$ similar examples as candidates. We utilize Sentence-BERT (SBERT) (Reimers and Gurevych, 2020) [^1] We use version distiluse-base-multilingual-cased-v1. for candidate selection. Based on our experiments detailed in Section B, we set $k=10$. The top $k$ candidates for $q_{i}$ are denoted as $D_{i}=\{d_{i,1},\cdots,d_{i,k}\}$, where each candidate $d_{i,j}$ is represented as $(x_{i,j},y_{i,j})$, with $x_{i,j}$ being the input and $y_{i,j}$ the corresponding label.
@@ 2.3 Retriever Fine-tuning
We utilize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023) to train the task-specific retriever, aiming to maximize the similarity between $q_{i}$ and $x_{i,j}$ if $x_{i,j}$ is a positive example while minimizing the similarity if $x_{i,j}$ is a negative example. We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant cost of fine-tuning an LLM. We train for 50 epochs using the AdamW optimizer with a learning rate of 2e-5 and a batch size of 16. Due to the multilingual nature of Glot500, the fine-tuned retriever can be effectively transferred to retrieve in-context examples for other languages.
@@ 3.1 Setup
We evaluate XAMPLER on two text classification benchmarks: SIB200 (Adelani et al., 2023a) and MasakhaNEWS (Adelani et al., 2023b). The partitioning for these datasets was predefined by the respective benchmarks. SIB200 is a massively multilingual text classification benchmark with seven classes. Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500 (see §C). The English training set contains 701 samples, with 204 evaluation samples per language. MasakhaNEWS is a news classification task for 16 African languages, covering six topics. The English training set contains 3.31k samples, with 175 to 948 evaluation samples per language.
Our evaluation framework follows the prompt template used in Lin et al. (2024b): ‘The topic of the news [sentence] is [label]’, where [sentence] represents the text for classification and [label] is the ground truth. [label] is included when the sample serves as a few-shot example but is omitted when predicting the sample. We opt for English prompt templates over in-language ones due to the labor-intensive nature of crafting templates for non-English languages, especially those with limited resources. MaLA500 takes the concatenation of few-shot examples and $q_{i}$ as input, then proceeds to estimate the probability distribution across the label set. We measure the performance with accuracy.
@@
We compare XAMPLER with the following retrieving strategies:
@@
Multilingual Language Models. We use two massively multilingual language models, Glot500 and MaLA500, as retrievers. Tailored examples are retrieved based on the cosine similarity between the sentence representations of the candidate and the query. For Glot500, we utilize mean pooling over hidden states of the selected layer. For MaLA500, we adopt a position-weighted mean pooling method on the selected layer, assigning higher weights to later tokens (Muennighoff, 2022). We use K-Nearest Neighbors to select the layer that performs best across layers (see §A). The selected layers for Glot500 and MaLA500 are 11 and 21, respectively.
Off-the-shelf Retriever. We also employ three off-the-shelf retrievers trained on parallel data: SBERT (Reimers and Gurevych, 2019), LaBSE (Feng et al., 2022), and Multilingual E5 (Wang et al., 2024). [^2] https://huggingface.co/intfloat/multilingual-e5-large
We set the number of shots as the number of classes and evaluate under two settings: the label-aware setting, where one shot is provided per class, and the label-agnostic setting, where the most similar examples are retrieved regardless of their labels.
SIB200 MasakhaNEWS
Label-Aware Label-Agnostic Label-Aware Label-Agnostic
Random 65.24 61.68 72.32 72.39
Glot500 66.60 68.55 73.35 73.01
MaLA500 66.75 66.25 73.39 71.58
SBERT 67.13 66.59 73.24 72.8
LaBSE 68.51 73.69 72.54 73.29
Multilingual E5 69.09 74.61 73.63 72.61
XAMPLER 70.18 75.91 75.02 73.85
Table 1: Average macro-accuracy across the evaluated languages on SIB200 and MasakhaNEWS using XAMPLER compared to the baselines.
@@ 3.2 Main Results
The comparison between the baselines and XAMPLER is illustrated in Table 1. Our analysis reveals several insights based on the performance with In-Context Learning (ICL) across different methods. Notably, the random baseline exhibits the worst performance among the baselines using ICL, emphasizing the critical role of example selection for effective in-context learning. Multilingual language models, which retrieve examples based on semantic similarity learned during pretraining, slightly outperform the random baseline. Leveraging an off-the-shelf retriever further improves performance, with Multilingual E5 emerging as the top performer among them. Among all the methods, XAMPLER achieves the highest performance in in-context learning. Specifically, on SIB200, XAMPLER surpasses Multilingual E5 by 1.09% in the label-aware setting and 1.30% in the label-agnostic setting, and by 1.93% and 1.24% on MasakhaNEWS, respectively.
Our two established practices to reduce resource requirements, i.e., selecting the top-k similar examples as candidates using existing off-the-shelf retrievers and using a smaller base model for retriever training, ensures that XAMPLER operates efficiently. For example, on the SIB200 benchmark (701 samples, 10 candidates per sample) using an NVIDIA GeForce GTX 1080 Ti (11GB), retriever fine-tuning took just 1.5 hours, and retrieving similar examples added only 0.1 seconds per query during inference.
3.3 Ablation Study
To further assess the effectiveness of XAMPLER, we conduct the following ablation studies:
Method Accuracy (%)
XLT (Glot500) 69.51
XLT (MaLA500) 69.90
MT 74.50
KNN 72.85
XAMPLER 75.91
Table 2: Performance comparison of XAMPLER with ablation methods on the SIB200 benchmark.
Cross-lingual Transfer (XLT). Cross-lingual transfer is another approach that utilizes English data. In this method, the multilingual language model is fine-tuned on English data and then evaluated across target languages. Both Glot500 and MaLA500 are included, with their respective cross-lingual transfer methods denoted as XLT (Glot500) and XLT (MaLA500). For Glot500, we perform full-parameter fine-tuning with a batch size of 16 and a learning rate of 1e-5. For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we update only the LoRA parameters with prompt tuning. The learning rate is set to 1e-3, weight decay is 0.1, the maximum sequence length is 128, and the batch size is 16. The optimizer used is AdamW.
Machine Translation (MT). We translate the examples retrieved by XAMPLER from English to the target language and use these translated examples as few-shot examples for in-context learning. For translation, we use the distilled 600M variant of NLLB-200 (Costa-jussà et al., 2022).
K-Nearest Neighbors (KNN). We consider KNN with the fine-tuned task-specific retriever of XAMPLER for comparison. Specifically, we adopt majority voting based on the labels of the examples retrieved by the given retriever.
@@ 3.3 Ablation Study
The results of the ablation studies are shown in Table 2. XAMPLER outperforms XLT (Glot500) and XLT (MaLA500) by 6.40% and 6.01%, respectively, demonstrating that in-context learning, when combined with informative cross-lingual few-shot examples, is superior to traditional cross-lingual transfer methods via fine-tuning. Translating English examples into target languages results in slightly worse results than XAMPLER by 1.41%. It demonstrates that XAMPLER benefits more from English in-context examples than in-language examples. A comparison between XAMPLER and KNN shows that XAMPLER performs better by 3.06%. We also compared XAMPLER’s performance with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 2. Interestingly, XAMPLER with ICL exhibits inconsistent superiority over KNN, with performance variances ranging from 3% to 10%. Specifically, XAMPLER with KNN achieves its peak performance with 5 examples, whereas ICL achieves impressive results with only 2 examples. Notably, in comparison to KNN’s optimal performance, recorded at 73.26% with 5 shots, XAMPLER with ICL demonstrates a notable improvement of 2.58%. These findings underscore the efficacy of applying in-context learning in effectively leveraging the retrieved examples.
@@ 4 Related Work
Several works (Winata et al., 2021; Shi et al., 2022; Winata et al., 2022; Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023; Cahyawijaya et al., 2024) extend these methods to non-English languages. A study closely related to ours is Shi et al. (2022), which trains a cross-lingual example retriever via distilling the LLM’s scoring function and evaluates it on four languages for the Text-to-SQL Semantic Parsing task. However, our contribution lies in addressing the more challenging low-resource scenario, thereby extending the applicability and robustness of the approach proposed by Shi et al. (2022).
@@ 5 Conclusion
In this paper, we introduce XAMPLER, a novel approach designed for cross-lingual example retrieval to facilitate in-context learning in any language. Relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language. Experiments on SIB200 and MasakhaNEWS show that XAMPLER outperforms previous methods by a notable margin.